OpenStudy (samigupta8):

Two blocks of masses m1=1kg and m2=2kg are connected by a non deformed light spring. They are lying on a rough horizontal table.The coefficient of friction between blocks and surface is 0.4.What minimum constant force F has to be applied in horizontal direction to the block of mass m1 in order to shift the other block?

OpenStudy (samigupta8):

@agent0smith

OpenStudy (anonymous):

|dw:1458012634121:dw The point at which block m2 moves will the static equilibrium of block m2 sum of Fx=0 to block m2 T-F2 = 0 T = F2 This suggests that T would be sufficient to move block m2 Sum of Fx=0 to block m1 F_H - T - F1 = 0 F_H = T + F1 T = F2 Horizontal force = F1 + F2

OpenStudy (anonymous):